Can you help me solve this problem: 1/2n + 6 = 1/3n
I have subtracted the 1/2n from the left side to combine with the right side. I have 6 = 5/6n. I'm stuck now. Just to clarify the / indicates a fraction.

Answer by stanbon(75887)

(1/2)n + 6 = (1/3)n
-------
Multiply thru by 6 to get rid of the fractions:
3n + 36 = 2n
n = -36
----------------------
Cheers,
Stan H.